OpenStudy (anonymous):

Write the given sentence as an equation. Tim’s age in 7 years will be three times what it was 19 years ago.

OpenStudy (anonymous):

these are the choices i have A. 3(t + 7) = t – 19 B. t + 19 = 3(t – 7) C. t + 7 = 3(t – 19) D. 3(t + 19) = t – 7

OpenStudy (paxpolaris):

what do you think it is?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

idk if im right ??

OpenStudy (paxpolaris):

break the sentence apart ... and write each part ...as part of equation

OpenStudy (paxpolaris):

Tim’s age in 7 years ... in 7 yrs tim's age will b :

OpenStudy (anonymous):

14 ?

OpenStudy (paxpolaris):

t+7

OpenStudy (anonymous):

so has to be C ??

OpenStudy (paxpolaris):

will be/is/was means =

OpenStudy (paxpolaris):

three times \[3\times(?)\] so far we have: \[\Large t+7=3\times \left( ? \right)\]

OpenStudy (paxpolaris):

3 times what : what it was 19 years ago t-19 so, yes C

OpenStudy (anonymous):

thank you cus i get confuse sometimes